Two factory plants are making TV panels. Yesterday, Plant A produced 3000 panels. Ten percent of the panels from Plant A and 3%of the panels from Plant B were defective. How many panels did Plant B produce, if the overall percentage of defective panels from the two plants was 6% ?

Answer by ikleyn(52781)

.
Two factory plants are making TV panels. Yesterday, Plant A produced 3000 panels.
Ten percent of the panels from Plant A and 3%of the panels from Plant B were defective.
How many panels did Plant B produce, if the overall percentage of defective panels from the two plants was 6% ?
~~~~~~~~~~~~~

Let x = How many panels did Plant B produce.

The number of defective panels from plant A is 0.1*3000.

The number of defective panels from plant B is 0.03*x


The total defective panels from both plants is

    0.1*3000 + 0.03x = 0.06*(3000+x).


Simplify and find x

    300 + 0.03x = 180 + 0.06x

    300 - 180 = 0.06x - 0.03x

       120    =     0.03x

        x     =     120/0.03 = 4000.


ANSWER.  Plant B produced 4000 panels.

Solved.